OpenStudy (anonymous):

Find the points on the curve y = 2x^3 + 3x^2 - 12x + 1 where the tangent is horizontal.

OpenStudy (anonymous):

take the derivative and set it equal to 0, find x then solve for y in the original equation

OpenStudy (anonymous):

y'=6x^2+6x-12

OpenStudy (anonymous):

6x^2+6x-12=0 x^2+x-2=0 (x+2)(x-1)=0 x=-2, 1

OpenStudy (anonymous):

plug these values of x into y = 2x^3 + 3x^2 - 12x + 1 to find the y-coordinate

OpenStudy (anonymous):

so (-2,0) and (1,0) are the answers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you plug x=-2 and x=1 into the original equation, not the derivative

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

so (-2,21) and (1,-6)

OpenStudy (anonymous):

right